OpenStudy (anonymous):

Disprove the following a=x a+a=a+x 2a=a+x 2a-2x = a+x-2x 2(a-x) = a+x-2x 2(a-x) = a-x 2=1 Problem?

OpenStudy (amistre64):

as my dad says: "Because I said so!!" :)

OpenStudy (anonymous):

divide by zero.

OpenStudy (anonymous):

dividing by zero in the step 2(a-x) = a+x-2x 2(a-x) = a-x

OpenStudy (anonymous):

well actually in this step 2(a-x) = a-x 2=1 left hand side is \[2\times 0\] right hand side is 0 cannot divide by zero to show that 2 = 1

OpenStudy (anonymous):

Haha well done